When we say "with respect to”. Like a with respect to b then what do we have to do? a-b or b-a or b/a please help . (While finding relative refractive index, etc) 
 

Grade:11

3 Answers

Umakant biswal
7 years ago
@ siddhi 
thats a good question . 
absolute refractive index can be measured only by ratio , of the two diffrent medium to be considered . 
for example – refractive index of medium 2 with respect to medium 1 will be given as 
u2/u1 = c/v2 / c/v1 = v1/v2 
so, obviously it will be a/b 
HOPE IT CLEARS NOW . 
ALL THE BEST ..
Siddhi
7 years ago
@ umakant biswal then in velocitys (of two trains when going in same direction or opp) why do we subtract or add. There also it is “with respect to"
 
Umakant biswal
7 years ago
@ siddhi 
that is relative velocity and here it is optical part , both are opposite of each other . 
and their we take that with respect to ground and ground is a frame of reference or stationary point . 
here we are taking 2 medium in concept and here no point works as frame of reference . 
HOPE IT CLEARS NOW ..